# Modular knots obey the Chebotarev law

### v3

@@ Abstract.
We refine McMullen’s construction of sequences of knots in $S^{3}$ obeying the Chebotarev law by invoking the notion of rational Fried surgeries, and improve argument for the planetary links. Furthermore, we point out that the set of modular knots around any torus knot in $S^{3}$ together with the missing torus knot also obey the Chebotarev law, and deduce an equidistribution theorem for the classical Rademacher symbol.
@@ 1. Introduction
A Chebotarev link in $S^{3}$ is an analogue of the set of all prime numbers in ${\mathbb{Z}}$, and would play an important roll in arithmetic topology, especially when we formulate an analogue of the idelic class field theory for 3-manifolds [Uek21] (see also [Mor12, Nii14, NU19, Mih19, KMNT21, NU23, Tas24]). Here is the definition:
Definition 1 ((The Chebotarev law)).
@@
A countably infinite link ${\mathcal{K}}$ is said to be Chebotarev if it obeys the Chebotarev law with respect to some order.
In order to answer Mazur’s question on the existence of a Chebotarev arrangement in [Maz12], McMullen proved a highly interesting theorem, generalizing Adachi–Sunada’s result [Sun84, Proposition II-2-12], by using Parry–Pollicott’s theory on zeta functions of symbolic dynamics [PP90]:
Proposition 2 ([McM13, Theorem 1.2]).
@@
Let $(K_{i})$ be the closed orbits of any topologically mixing pseudo-Anosov flow on a closed 3-manifold $M$, ordered by length in a generic metric. Then $(K_{i})_{i}$ obeys the Chebotarev law. (We consult [Mos92, Fen08, Cal07] for the terminology of pseudo-Anosov flows.)
By applying his theorem to the monodromy suspension flow of the figure-eight knot $K$ and noting that the Chebotarev law persists under Dehn surgeries, he constructed a Chebotarev link containing $K$ in $S^{3}$ [McM13, Corollary 1.3], claiming that his construction is applicable to any hyperbolic fibered knot. In this article, we refine his construction to verify his claim for some subtle cases (see Section 2), as well as to improve the author’s argument in the previous paper [Uek21, Propositions 12, 15]. Namely, we first aim to verify the following assertions.
Theorem 3.
@@
Let $L$ be a fibered hyperbolic link in $S^{3}$ and let $(K_{i})$ denote the sequence of knots consisting of the closed orbits of the suspension flow of the monodromy map and $L$ itself. Then $(K_{i})$ obeys the Chebotarev law if ordered by length with respect to a generic metric.
@@
The union ${\mathcal{L}}=\cup_{i}K_{i}$ is a stably Chebotarev link, that is, for any finite branched cover $h:M\to S^{3}$ branched along any finite link in ${\mathcal{L}}$, the inverse image $h^{-1}({\mathcal{L}})$ is again Chebotarev.
@@
For this purpose, we introduce the notion of a generalized pseudo-Anosov flows, which allows 1-pronged singular orbits (Section 3), invoke rational Fried surgeries (Section 4), and establish a version of 2 for such flows (7).
@@
We remark that the Chebotarev link ${\mathcal{L}}$ in 3 is also called the planetary link rising from $L$ and sometimes admits another amazing property due to Ghrist and others; the link ${\mathcal{L}}$ may contain every type of links. In Section 5, we prove 3 and attach a fused question.
We also provide new examples of Chebotarev links:
Theorem 4.
@@
Let $(a,b)$ be any pair of coprime integers. Then, the set of modular knots around the $(a,b)$-torus knot $K_{a,b}$ in $S^{3}$ and $K_{a,b}$ obeys the Chebotarev law if ordered by length in a generic metric.
Here, a modular knot is defined to be a closed orbit of the so-called geodesic flow on the unit tangent bundle of the orbifold $\Gamma_{a,b}\setminus{\mathbb{H}}$ of the $(a,b,\infty)$-triangle group or its certain finite cover homeomorphic to $S^{3}\setminus K_{a,b}$. As a corollary, we obtain an equidistribution formula for a certain ubiquitous function called the Rademacher symbol $\Psi:\mathop{\mathrm{SL}}\nolimits_{2}{\mathbb{Z}}\to{\mathbb{Z}}$ (Corollary 9). We present the details in Sections 68.
In Section 9, we remark further interests of Chebotarev links, such as the Lang–Trotter conjecture and the Neukirch–Uchida theorem, from a viewpoint of arithmetic topology.
@@ 2. McMullen’s construction and Dehn surgeries
Let us recall McMullen’s construction; The mapping torus of $A={\small\begin{pmatrix}1&1\\ 1&2\end{pmatrix}}$ with $\mathop{\mathrm{tr}}\nolimits A>2$ acting on $T={\mathbb{R}}^{2}/{\mathbb{Z}}^{2}$ is the result $K(0)$ of the 0-surgery along the figure-eight knot $K=4_{1}$ in $S^{3}$ [BZH14, Chapter 5], so that the monodromy suspension flow is a pseudo-Anosov flow on the torus bundle $K(0)$ such that the exterior of the 00-orbit is homeomorphic to $S^{3}-K$. By [McM13, Corollary 2.2], the flow is topologically mixing if there are two orbits $K_{i}$ and $K_{j}$ whose length satisfy $\ell(K_{i})/\ell(K_{j})\not\in{\mathbb{Q}}$, hence a generic time change makes the pseudo-Anosov flow to be topologically mixing. Therefore by [McM13, Theorem 1.2], the closed orbits in $K(0)$ obey the Chebotarev law. Since the Chebotarev law persists under Dehn surgeries along knots, the $\infty$-surgery along the 00-orbit yields a sequence of knots in $S^{3}$ containing $K$ itself and obeying the Chebotarev law.
@@
Indeed, Gabai [Gab97, p.27] initially pointed out by using Penner’s program that the 0-surgery along $K=8_{20}$ results in a fibered manifold with reducible monodromy, so that the mapping torus is toroidal, hence not hyperbolic. This fact is known as the consequence of the fact that the complement of $8_{20}$ had an essential punctured torus with boundary slope 0 [HO89, p.479]. Here for the convenience of readers, we give a slightly generalized assertion (Proposition 6) suggested by Motegi. The crosscap number of a knot in the 3-sphere is defined as the minimal first Betti number of non-orientable subsurfaces bounded by the knot. Note that $K=8_{20}$ is the pretzel knot $P(2,3,-3)$ and its crosscap number is two.
Proposition 6.
@@
Consider the decomposition $K(0)=X\cup Y$ with $X\cap Y=T$. If $T$ is incompressible in $Y$ as well, then so it is in $K(0)$, and hence $K(0)$ is toroidal. Let us suppose that $T$ is compressible in $Y$ to deduce contradiction; If $Y$ is irreducible, then $Y$ is a solid torus, hence $K(0)$ is a prism manifold or ${\mathbb{R}}{\rm P}^{3}\#{\mathbb{R}}{\rm P}^{3}$. In the former case, $\pi_{1}(K(0))$ is a finite group, contradicting $H_{1}(K(0))\cong{\mathbb{Z}}$. In the latter case, $K(0)$ is reducible, contradicting Gabai’s great result [Gab87, Corollary 8.3]. If $Y$ is reducible, then $Y$ is the connected sum $V\#Z$ of a solid torus $V$ and some closed 3-manifold $Z\neq S^{3}$. Hence we have $K(0)=W\#Y$ for a prism manifold $W$ or $K(0)={\mathbb{R}}{\rm P}^{3}\#{\mathbb{R}}{\rm P}^{3}\#Z$, again contradicting Gabai’s result. ∎
@@
In addition, hyperbolic surgeries do not always result Seifert fibered spaces, so that we do not necessarily obtain a pseudo-Anosov flow on a $S^{1}$-bundle over a closed surface.
3. Generalized pseudo-Anosov flows
@@
One may expect the following: Since the whole argument for pseudo-Anosov flows in [Col79] applies, a generalized pseudo-Anosov flow still admits a Markov section, so that the proof of [McM13, Theorem 1.2] (contained in Proof of Theorem 1.4) works as well, yielding the following assertion.
Theorem 7.
@@
Instead of precisely verifying the argument above, we will verify this assertion by invoking the notion of rational Fried surgeries in the next section.
4. Rational Fried surgeries
Let $K$ be a closed orbit of an Anosov flow and $\sigma$ a cross section in the boundary of the tubular neighborhood of $K$ such that $\sigma$ transversely intersects each of the stable/unstable flow lines twice. Then a Fried surgery along $(K,\sigma)$ naturally induces an Anosov flow on the result of the Dehn surgery whose slope coincides with that of $\sigma$. It is defined as the combination of the blow-up along $K$ and the blow-down by parallel copies of $\sigma$, so that the orbits of the flow and the stable/unstable foliations persist in the exterior [Fri83]. We remark that Fried surgeries are topologically equivalent to Goodman surgeries [Goo83, Sha20].
Proof of 7.
If we instead apply a similar procedure for a closed orbit of a generalized pseudo-Anosov flow and a cross section $\sigma$ such that $\sigma$ transversely intersects each of the stable/unstable flow lines more than once, then we mostly obtain a $k$-pronged orbit with $k\geq 2$ and rarely a 1-pronged orbit in a new flow on the result of the corresponding rational Dehn surgery. Note in addition that all but finite rational surgeries are hyperbolic by Thurston [Thu82].
Therefore, we obtain the assertion by combining the 00-fillings, rational Fried surgeries, a generic time change, [McM13, Theorem 1.2], the $\infty$-Dehn surgeries, and Proposition 5. ∎
@@ 5. The planetary links
We prove Theorem 3 by using 7 and 5:
Proof of Theorem 3.
@@
A pseudo-Anosov map on a punctured surface $\Sigma$ extends to a map on a closed surface ${\widehat{\Sigma}}$ such that the stable/unstable foliations may admit 1-pronged singular closed orbits [Thu88, CB88, FM12]. The pseudo-Anosov suspension flow on the mapping torus of $\Sigma$ extends to a generalized pseudo-Anosov flow on that of ${\widehat{\Sigma}}$ such that the components of $L$ are contained in the set of the closed orbits. By [McM13, Corollary 2.2] based on [PP90, Theorem 8.5], a generic time change makes the flow topologically mixing. Hence Theorem 7 yields a sequence of knots obeying the Chebotarev law.
Applying Proposition 5 to the $\infty$-Dehn surgeries along $L$ in the mapping torus of the closed surface ${\widehat{\Sigma}}$, we obtain a sequence of knots in $S^{3}$ obeying the Chebotarev law, proving the first assertion.
Since a generalized pseudo-Anosov flow lifts to that on a finite branched cover branched along a finite set of closed orbits, we clearly obtain the second assertion. ∎
@@
The Chebotarev link ${\mathcal{L}}=\cup_{i}K_{i}$ obtained by Theorem 3 is called the planetary link rising from $L$, after Birman–Williams [BW83]. Ghrist and others proved that if $L$ belongs to a certain large class of links (eg, the figure-eight knot, the Whitehead link, the Borromean ring, and every fibered non-torus 2-bridge knot), then the planetary link ${\mathcal{L}}$ contains all types of links [Ghr97, GHS97, GK04].
@@
In order to answer his fused question in the affirmative, it suffices to show the equidistribution of each type of knot in the planetary link. Miller’s work [Mil01] might give a cliff, which explicitly finds some types of knots in the planetary link. We leave this question for a future study.

### v5

@@ Abstract.
We study knots which behave like prime numbers. We discuss the planetary link raised from a hyperbolic fibered link in $S^{3}$ with an emphasis on surgeries, point out certain subtleness, and refine the construction. In addition, we point out a version of McMullen’s theorem for the cases over cusped orbifolds and deduce that the family of modular knots around any torus knot $K_{a,b}$ in $S^{3}$ together with the missing knot $K_{a,b}$ obey the Chebotarev law. Furthermore, we attach several remarks in the view of arithmetic topology.
@@ 1. Introduction
The analogy between knots and primes has played important roles from the era of Gauss. In modern times, the analogy was initially pointed out by Mazur [Maz64] with an emphasis on Iwasawa theory and Alexander–Fox theory, and has been systematically and vastly developed by Kapranov, Reznikov, Morishita, Kim, and others [Kap95, Rez97, Rez00, Mor02, Mor12, Kim20]. This research field is called arithmetic topology. Mazur [Maz12] further asked if there exists a sequence of knots in $S^{3}$ that behave like the set of all prime numbers in ${\mathbb{Z}}$, namely, that obeys the Chebotrev law. Here is the definition:
Definition 1.1 ((The Chebotarev law)).
@@
In this view, McMullen proved the following highly interesting theorem, generalizing Adachi–Sunada’s result [Sun84, Proposition II-2-12].
Theorem 1.2 ([McM13, Theorem 1.2]).
@@
Let $(K_{i})$ be the closed orbits of any topologically mixing pseudo-Anosov flow on a closed 3-manifold $M$, ordered by length in a generic metric. Then $(K_{i})_{i}$ obeys the Chebotarev law.
We consult [Mos92, Fen08, Cal07] for the terminology of pseudo-Anosov flows. A key of the proof is the following: since such flows admit Markov partitions [FS79], the flow dynamics admit symbolic encodings, so the theory of symbolic dynamics (cf. [PP90]) applies.
By using this theorem, McMullen pointed out that the family ${\mathcal{L}}$ of closed orbits of the monodromy suspension flow of the figure-eight knot $K=4_{1}$ together with $K$ in $S^{3}$ obeys the Chebotarev law if ordered by length [McM13, Corollary 1.3], answering Mazur’s question. He also claimed that the same construction works for any hyperbolic fibered knot in $S^{3}$.
However, there is something subtle to care about, even in the case of $4_{1}$. In the former half of this article, we aim to verify his claim, as well as to improve the author’s argument in the previous paper [Uek21, Propositions 12, 15]. Namely, we re-prove the following assertion.
Theorem 1.3.
@@
(1) Let $L$ be a fibered hyperbolic finite link in $S^{3}$ and let ${\mathcal{L}}$ denote the union of the closed orbits of the suspension flow of the monodromy map and $L$ itself. Then the components of ${\mathcal{L}}$ obey the Chebotarev law if ordered by length in a generic metric.
@@
(2) For any finite branched cover $h:M\to S^{3}$ branched along a finite sublink of ${\mathcal{L}}$, the inverse image $h^{-1}({\mathcal{L}})$ obey the Chebotarev law, if ordered by length in a generic.
@@
For this purpose, in Section 2, we first recall his construction with an emphasis on Dehn surgeries and observe how subtle things may occur. Afterward, in Section 3, we introduce the notion of a generalized pseudo-Anosov flows, which allows 1-pronged singular orbits, invoke rational Fried surgeries, and establish a version of 1.2 for such flows (3.1). Finally, we prove 1.3 in Section 4.
@@
We remark that the Chebotarev link ${\mathcal{L}}$ in 1.3 is also called the planetary link raised from $L$ and sometimes admits another amazing property due to Ghrist and others [Ghr97, GHS97, GK04]; this link ${\mathcal{L}}$ contains every type of links. We also attach a fused question.
The modular knots around the trefoil $K_{2,3}=3_{1}$ in $S^{3}$ is another famous example of infinite links of deep features, with a close connection to Lorenz knots (cf. [BW83, Tuc99, Ghy07], see Section 5 for more details). Since the planetary link raised from $4_{1}$ and the family of modular knots are like siblings, it is desirable to have a similar result for modular knots. However, each of them has its specific nature and requires slightly different arguments. In this view, we would like to point out the following partial extension of 1.2.
Theorem 1.4.
(1) Let $\Gamma\leq{\rm SL}_{2}{\mathbb{R}}$ is any Fuchsian group, so that the orbifold $\Σ\Γ\Sigma=\Gamma\backslash{\mathbb{H}}$ may have cusps. Let $L$ be a finite link in a closed 3-manifold $M$ and suppose that the exterior $M-L$ is the unit tangent bundle $T_{1}\Sigma$. Then, the closed orbits of (a lift of) the geodesic flow and the missing link $L$ obey the Chebotarev law, if ordered by length in a generic metric.
(2) If instead $M-L$ is a finite cover of $T_{1}\Sigma$, then a similar result holds.
Now let $(a,b)$ be any pair of coprime integers and let $\Gamma_{a,b}\leq{\rm SL}_{2}{\mathbb{R}}$ denote the $(a,b,\infty)$-triangle group. Then a modular knot is defined to be a closed orbit of a standard flow called the geodesic flow on (i) the unit tangent bundle of the orbifold $\Gamma_{a,b}\backslash{\mathbb{H}}$, which is homeomorphic to the exterior $L(a,b)-\overline{K}_{a,b}$ of a certain knot in the lens space, or (ii) its certain finite cover homeomorphic to $S^{3}-K_{a,b}$, where $K_{a,b}$ is the $(a,b)$-torus knot. More precisely, we have $\pi_{1}L(a,b)={\mathbb{Z}}/r{\mathbb{Z}}$ with $r=ab-a-b$ and the corresponding ${\mathbb{Z}}/r{\mathbb{Z}}$-cover is $S^{3}-K_{a,b}\to L(a,b)-\overline{K}_{a,b}$ (cf. [Tsa13]). The classical case is $\Gamma_{2,3}={\rm SL}_{2}{\mathbb{Z}}$, $K_{2,3}=$ (trefoil), and $r=1$.
1.4 particularly yields the following.
Theorem 1.5.
@@
For any pair $(a,b)$ of coprime integers, (i) the family of modular knots in $L(a,b)-\overline{K}_{a,b}$ and the missing knot $\overline{K}_{a,b}$ and (ii) the family of modular knots in $S^{3}-K_{a,b}$ and the missing torus knot $K_{a,b}$ obey the Chebotarev law, if ordered by length in a generic metric.
As a corollary, we obtain an equidistribution formula for a certain ubiquitous function called the Rademacher symbol $\Psi:\Gamma_{2,3}=\mathop{\mathrm{SL}}\nolimits_{2}{\mathbb{Z}}\to{\mathbb{Z}}$ (Corollary 6.1). This may be seen in contrast with Sarnak–Mozochi’s another distribution theorem. See Section 6 for the details.
The idelic class field theory for number fields developed by Artin–Takagi and Chevalley sums up all local theories associated to primes to control the global abelian extensions with ramifications. Its topological analogue for a 3-manifold endowed with a nice infinite link has been developed in the series of works [Nii14, NU19, Mih19, KMNT21, NU23, Tas24, Tas25]. As proved in [Uek21] (see also the latest survey [Mor24, Section 7]), a link that obeys the Chebotarev law is stably generic in the sense of Mihara [Mih19] and guarantees a certain functionality. In Section 7, we remark further possible interests of Chebotarev links, such as the Lang–Trotter conjecture and the Neukirch–Uchida theorem, from the viewpoint of arithmetic topology.
@@ 2. McMullen’s construction and Dehn surgeries
Let us recall McMullen’s construction. The mapping torus of $A={\small\begin{pmatrix}1&1\\ 1&2\end{pmatrix}}$ with $\mathop{\mathrm{tr}}\nolimits A>2$ acting on $T={\mathbb{R}}^{2}/{\mathbb{Z}}^{2}$ is the result $K(0)$ of the 0-surgery along the figure-eight knot $K=4_{1}$ in $S^{3}$ [BZH14, Chapter 5], so that the monodromy suspension flow is a pseudo-Anosov flow on the torus bundle $K(0)$ such that the exterior of the 00-orbit is homeomorphic to $S^{3}-K$. By [McM13, Corollary 2.2], the flow is topologically mixing if there are two orbits $K_{i}$ and $K_{j}$ whose length satisfy $\ell(K_{i})/\ell(K_{j})\not\in{\mathbb{Q}}$, hence a generic time change makes the pseudo-Anosov flow to be topologically mixing. Therefore by [McM13, Theorem 1.2], the closed orbits in $K(0)$ obey the Chebotarev law. Since the Chebotarev law persists under Dehn surgeries along knots, the $\infty$-surgery along the 00-orbit yields a sequence of knots in $S^{3}$ containing $K$ itself and obeying the Chebotarev law.
@@
Indeed, Gabai [Gab97, p.27] initially pointed out by using Penner’s program that the 0-surgery along $K=8_{20}$ results in a fibered manifold with reducible monodromy, so that the mapping torus is toroidal, hence not hyperbolic. This fact is known as the consequence of the fact that the complement of $8_{20}$ had an essential punctured torus with boundary slope 0 [HO89, p.479]. Here for the convenience of readers, we give a slightly generalized assertion (Proposition 6) suggested by Motegi. The crosscap number of a knot in the 3-sphere is defined as the minimal first Betti number of non-orientable subsurfaces bounded by the knot. Note that $K=8_{20}$ is the pretzel knot $P(2,3,-3)$ and its crosscap number is two [IM10, Theorem 1.2].
Proposition 2.2.
@@
Consider the decomposition $K(0)=X\cup Y$ with $X\cap Y=T$. If $T$ is incompressible in $Y$ as well, then so it is in $K(0)$, and hence $K(0)$ is toroidal. Let us suppose that $T$ is compressible in $Y$ to deduce contradictions; If $Y$ is irreducible, then $Y$ is a solid torus, hence $K(0)$ is a prism manifold or ${\mathbb{R}}{\rm P}^{3}\#{\mathbb{R}}{\rm P}^{3}$. In the former case, $\pi_{1}(K(0))$ is a finite group, contradicting $H_{1}(K(0))\cong{\mathbb{Z}}$. In the latter case, $K(0)$ is reducible, contradicting Gabai’s great result [Gab87, Corollary 8.3] asserting that the result of the zero surgery along a knot in $S^{3}$ is a prime manifold. If $Y$ is reducible, then $Y$ is the connected sum $V\#Z$ of a solid torus $V$ and some closed 3-manifold $Z\neq S^{3}$. Hence we have $K(0)=W\#Y$ for a prism manifold $W$ or $K(0)={\mathbb{R}}{\rm P}^{3}\#{\mathbb{R}}{\rm P}^{3}\#Z$, again contradicting Gabai’s result. ∎
@@
3. Generalized pseudo-Anosov flows and rational Fried surgeries
@@
One might expect the following: The whole argument for pseudo-Anosov flows in [FP79] persists, a generalized pseudo-Anosov flow admits a version of Markov partition, the proof of [McM13, Theorem 1.2] (contained in Proof of [McM13, Theorem 1.4]) works as well, and we have a version of 1.2 for a generalized pseudo-Anosov flow. Instead of precisely verifying this argument, we invoke the notion of rational Fried surgeries to prove the assertion.
We first recollect the classical case. Let $K$ be a closed orbit of an Anosov flow and $\sigma$ a cross section of $K$, namely, a circle on the boundary of the tubular neighborhood of $K$, that transversely intersects each of the stable/unstable flow lines twice. Then a Fried surgery along $(K,\sigma)$ naturally induces an Anosov flow on the result of the Dehn surgery whose slope coincides with that of $\sigma$. It is defined as the combination of the blow-up along $K$ and the blow-down by parallel copies of $\sigma$, so that the orbits of the flow and the stable/unstable foliations persist in the exterior [Fri83]. We remark that Fried surgeries are topologically equivalent to Goodman surgeries [Goo83, Sha20].
Next, we demonstrate a version of 1.2 for a generalized pseudo-Anosov flow:
Theorem 3.1.
@@
Proof of 3.1.
If we instead apply a similar procedure for a closed orbit $K$ of a generalized pseudo-Anosov flow and a cross section $\sigma$ of $K$ such that $\sigma$ transversely intersects each of the stable/unstable flow lines more than once, then we mostly obtain a $k$-pronged orbit with $k\geq 2$ and rarely a 1-pronged orbit in a new flow on the result of the corresponding rational Dehn surgery. Therefore, by applying the 00-fillings, rational Fried surgeries, a generic time change, [McM13, Theorem 1.2], the $\infty$-Dehn surgeries, and Proposition 2.1 in this order, then we obtain the assertion. ∎
@@ 4. The planetary links
We prove Theorem 1.3 by using 3.1 and 2.1:
Proof of 1.3.
@@
The pseudo-Anosov map on a punctured surface $\Sigma$ extends to a map on a closed surface ${\widehat{\Sigma}}$ such that the stable/unstable foliations may admit 1-pronged singular closed orbits [Thu88, CB88, FM12]. The pseudo-Anosov suspension flow on the mapping torus of $\Sigma$ extends to a generalized pseudo-Anosov flow on that of ${\widehat{\Sigma}}$ such that the components of $L$ are contained in the set of the closed orbits. By [McM13, Corollary 2.2] based on [PP90, Theorem 8.5], a generic time change makes the flow topologically mixing. Hence Theorem 3.1 yields a sequence of knots obeying the Chebotarev law. Applying Proposition 2.1 to the $\infty$-Dehn surgeries along $L$ in the mapping torus of the closed surface ${\widehat{\Sigma}}$, we obtain the assertion. ∎
@@
The Chebotarev link ${\mathcal{L}}=\cup_{i}K_{i}$ obtained by Theorem 1.3 is called the planetary link rising from $L$, after Birman–Williams [BW83]. Ghrist and others proved that if $L$ belongs to a certain large class of links (eg, the figure-eight knot, the Borromean ring, and every fibered non-torus 2-bridge knot), then the planetary link ${\mathcal{L}}$ contains all types of links [Ghr97, GHS97, GK04]. Ghrist [Ghr97, Section 5] claims that the Whitehead link also belongs to this class, and it is suggested as a reader’s exercise in [GHS97, Remark 3.2.20].
@@
To answer his fused question in the affirmative, it suffices to show the equidistribution of each type of knot in the planetary link. Miller’s work [Mil01] could give a cliff, which explicitly finds some types of geodesic knots.